### v1

# GUI Based Fuzzy Logic and Spatial Statistics for Unsupervised Microscopy Segmentation

Abstract
Brightfield microscopy imaging of unstained live cells remains a persistent challenge due to low contrast, temporal changes in specimen phenotypes, irregular illumination, and the absence of training labels. While deep learning (DL) methods (e.g., Cellpose 3.0) achieve state-of-the-art (SOTA) performance, they require extensive labeled data and heavy computational resources, and they often fail under uneven illumination. We present the first unsupervised segmentation framework combining spatial standard deviation from local mean (SSDLM), fuzzy logic, adjusted variograms, Moran’s I, and cumulative squared shift of nodal intensity (CSSNI) to address these limitations. Unlike deep learning models, our approach requires no annotations or retraining and operates through a user-friendly GUI tailored for non-programming users. The robustness and generality were validated on three datasets, including cross-domain data. We benchmark our method against 2023–2024 SOTA models, including Cellpose 3.0 and StarDist, using a dataset of unstained myoblast images. Our method achieves a significant improvement in segmentation performance, with an IoU increase of up to 48% and statistically validated superiority ($p<0.01$, Wilcoxon signed-rank test). Expert evaluation from two biologists further supports the segmentation quality (Cohen’s $\kappa>0.75$). The proposed algorithm is lightweight, interpretable, and computationally efficient, offering a practical and effective alternative for cell segmentation in label-free microscopy. The code, the dataset and the results are available for reproducibility [^1] The article is under review process. Code or Dataset may be provided upon request..
keywords:
Computer vision , image analysis , live cell imaging , quantitative microscopy , unsupervised segmentation
journal: CMIG
1 Introduction
The use of bright-field microscopy is ubiquitous in research and clinical diagnostics for observing specimens, routine analysis with different objectives [1]. The study of biological samples, accompanied by machine learning and data science [2], often suffers from significant challenges for studying unstained (naturally pigmented) specimen which appear as low-contrast & noisy objects under microscope.
Deep learning (DL) tools (e.g.Cellpose 3.0) [3] achieve impressive results for stained (high contrast or pigmented) labeled data, but they consistently fail under the low-contrast conditions characteristic of unstained samples due to their dependency on training data and sensitivity to noise.In addition, scarcity of sufficient annotated data hinders the DL training process. Therefore, segmenting unstained live cells remains a significant challenge involving faster, annotation-free and training-free analysis of live cells without killing, staining, or fixing them.
Our work bridges this critical gap by introducing a novel unsupervised methodology for live unstained cell culture images. The proposed methodology requires no annotations or pretraining while outperforming state-of-the-art (SOTA) approaches in challenging bright-field microscopy conditions.
The fundamental challenge in unstained bright-field microscopy lies in the weak contrast between cells and background, arising from subtle differences in light absorption and scattering [1]. Recent DL-based segmentation tools [3, 4] typically rely on strong contrast, limiting their effectiveness for live-cell studies where staining is impractical. This limitation becomes particularly acute when examining delicate specimens like primary cell cultures or patient-derived samples.
Challenges in Current Methods
Our analysis identifies that using ML/DL faces ten core challenges of existing approaches: (I) Low contrast & noise – Current SOTA models achieve IoU scores below 0.3 on unstained data (see §V), as they cannot distinguish cells from background noise without strong contrast cues. (II) Uneven illumination – Inconsistent brightness across the field of view (FOV) misleads ML/DL models. (III) Optical aberrations – Dust, air bubbles, glare, and slide imperfections introduce artifacts. (IV) Overlapping structures – Cells and tissues often overlap which complicates segmentation. (V) Annotation dependency – Microscopy data demands accurate annotation, requires expert knowledge, and is both error-prone and time-consuming. Another problem is scarcity of plenty of data. (VI) Data imbalance – Rare biological events (e.g., abnormal cells) lead to biased models that perform poorly on rare but critical features. (VII) Lack of standardization – Models trained on one dataset often fail to generalize to unseen datasets or imaging conditions due to variations in microscopes, lighting, and deviation from the sample preparation. (VIII) 2D limitations – Bright-field microscopy lacks depth information, obscuring 3D features. (IX) Training dependency – Supervised tools require extensive annotations [2], making them impractical for many live-cell applications. (X) Generalization gaps – Models trained on one dataset often fail under new imaging conditions [5], especially with varying illumination or optical setups.
Traditional solutions like spatial masking [6] or spectral filtering [7] either require physical modifications to the imaging system or fail to adapt to biological variability. Computational approaches like region growing [8] depend heavily on seed selection and distance metrics, making them unreliable for heterogeneous samples.
Our Approach
We present a fundamentally new framework that is found to be robust under the following assumptions: I) the environment involving optical conditions and the imaging system of bright-field microscopy is set up to achieve a nearly homogeneous image plane (defined in section §II.B); II) one of two image sets (either background or specimens) must have optical heterogeneity caused by its inherent texture. The propsed model combines:
  • 1.
    Spatial Standard Deviation from Local Mean (SSDLM): A novel statistical metric that works for measuring the homogeneity of the image set (Theorem 1, §II.B).
  • 2.
    Fuzzy logic system: GUI-controlled, facilitated dynamic system to resolve classification ambiguity.
  • 3.
    Cumulative Squared Shift of Nodal Intensity (CSSNI): A novel spatial statistic that quantifies local intensity variations more robustly than traditional gradient-based measures in low-contrast regimes (§II.C).
  • 4.
    Adjusted variogram analysis: Robust to uneven illumination optical artifacts.
Contributions
This work introduces the first fully unsupervised segmentation framework that consistently outperforms state-of-the-art (SOTA) models from 2023–2024 on unstained bright-field microscopy images. Our method achieves a 48% improvement in Intersection over Union (IoU) compared to Cellpose 3.0 (Table 3), with statistically significant gains ($p<0.01$, Wilcoxon signed-rank test), all while operating on standard CPU hardware.
The core contributions of this study are summarized as follows:
  • 1.
    Unsupervised segmentation with ethical advantage: Our method eliminates the need for staining or labeling, preserving cell viability and avoiding cytotoxic effects—critical for live-cell and longitudinal studies.
  • 2.
    Practicality and accessibility: The algorithm requires no training, annotations, or GPU resources, making it deployable in resource-limited biological labs and usable by non-programmers.
  • 3.
    Biological relevance: Working directly on label-free images, the method enables morphology analysis without experimental artifacts, capturing cells in their native physiological states.
  • 4.
    Software and reproducibility: The full implementation is released as open-source software with an intuitive, GUI-based interface tailored for biomedical researchers.
  • 5.
    Real-world validation: The framework is validated on challenging datasets, including primary cell cultures and clinical microscopy samples, demonstrating robust performance in diverse imaging conditions.
2 Academic Discussion & Background
2.1 Introducing Image Plane
Image Plane: A theoretical and physical 2DD space, which is a collection of pixels that are arranged in an array and have intensity values $I$ is called an image plane [9, 10, 11]. The intensity value for any pixel is determined as the outcomes of transmittance or reflectance function $T$, the illumination factor is governed by the physical plane of the sensor $L$ and optical distortions/aberrations or sensor non-uniformity $\Delta$ [12, 11] and hence $I$ encodes a convolution of optical, spatial, and noise characteristics that must be considered for accurate segmentation or enhancement [13]. Let $I^{\prime}(x^{\prime},y^{\prime})$ be the intensity function defined as:
$I=I^{\prime}(x^{\prime},y^{\prime})=T\left(P^{-1}(x^{\prime},y^{\prime},1)^{\top}\right)\cdot L(x^{\prime},y^{\prime})+\Delta(x^{\prime},y^{\prime}),$ (1)
where: $I^{\prime}(x^{\prime},y^{\prime})$ is the intensity function at coordinates $(x^{\prime},y^{\prime})$, $T$ is a 3D property (e.g., texture or reflectance), $P^{-1}$ is the inverse projection matrix mapping 2D image coordinates to 3D space, $L(x^{\prime},y^{\prime})$ is a 2D modulation term (e.g., lighting), $\Delta(x^{\prime},y^{\prime})$ represents noise or offsets, $I$ is the scalar intensity value at a specific point $(x^{\prime},y^{\prime})$.
2.2 Notion of Homogeneous Image Plane & Homogeneity
2.2.1 Homogeneous Image Plane
According to the hypothesis considered, in our context of microscopy, a homogeneous image plane is defined as an image plane where the equation (1) can be written as:
$\begin{split}I^{\prime}(x^{\prime},y^{\prime})=T(P^{-1}(x^{\prime},y^{\prime},1)^{T})\cdot L_{0}\end{split}$ (2)
Here, $L_{0}$ is a constant (or nearly constant) representing uniform illumination across the field of view (FOV). Any observed intensity variation arises solely from the object’s texture, not from inconsistencies in the optical system, illumination, or imaging medium. Nullifying the $\Delta$ terms corresponds to an idealization of the optical properties—i.e., aberrations, distortions, or other system-induced inconsistencies (such as those introduced by lenses or other optical elements) are assumed to be zero or negligible.
According to our hypothesis in microscopy, the background of an image should be homogeneous, as it lacks inherent texture. However, in reality, achieving a perfectly homogeneous plane is impossible, since neither $\Delta$ equals zero nor $L$ remains strictly constant. Therefore, we introduce a new computationally efficient metric (compared to traditional homogeneity measures based on the gray-level co-occurrence matrix (GLCM)) to measure homogeneity and establish a threshold for it.
2.2.2 Homogeneity of Region of Interest (ROI) in Image
Traditionally, homogeneity is measured with the help of gray level co-occurrence matrix (GLCM) [14] and formulated as:
$H_{d}^{\theta}=\sum_{i=1}^{n}\sum_{j=1}^{n}\frac{p_{d}^{\theta}(i,j)}{1+|i-j|}$
where $n$ is the total number of gray levels in the neighbourhood selected, $p_{d}^{\theta}(i,j)$ denotes the probability of a pixel pair having intensities $i$ and $j$ respectively at a certain $d$ distance and angle $\theta$. Calculating homogeneity is computationally expensive [14].
We measure mean and standard deviation of spatial standard deviation with respect to localized mean (SD of SSDLM) for a neighbourhood. Unlike the homogeneity measures in the GLCM, this technique has limitations in capturing the homogeneity along every sense of direction, but it is proved below that the SSDLM inversely varies with the metric homogeneity (if it exists) along any sense of direction.
Let us consider a $3\times 3$ neighbourhood around any pixel intensity $\omega_{(i,j)}$. In this case, $i$ and $j$ represent the cartesian coordinate of that pixel and $\omega$ is a function that maps to the intensity of that pixel.
$\begin{pmatrix}\omega_{(i-1,j-1)}&\omega_{(i-1,j)}&\omega_{(i-1,j+1)}\\ \omega_{(i,j-1)}&\omega_{(i,j)}&\omega_{(i,j+1)}\\ \omega_{(i+1,j-1)}&\omega_{(i+1,j)}&\omega_{(i+1,j+1)}\end{pmatrix}$
Therefore, considering $\overline{\omega}$ as the mean value of $\omega$, SSDLM for the neighbourhood of $\omega_{i,j}$ will be represented as:
$SSDLM_{(i,j)}=\sqrt{\sum_{i}\sum_{j}\frac{(\omega_{(i,j)}.-\overline{\omega})^{2}}{N}}$ (3)
From the context of GLCM matrix, we know that the homogeneity directly varies with the contrast [14]. Hence,
$\sum_{i=1}^{n}\sum_{j=1}^{n}\frac{p(i,j)}{1+|i-j|}\propto\frac{1}{\sum_{i=1}^{n}\sum_{j=1}^{n}(i-j)^{2}\cdot p(i,j)}$
$Again,\sum_{i=1}^{n}\sum_{j=1}^{n}(i-j)^{2}\cdot p(i,j)\propto\sqrt{\sum_{i}\sum_{j}\frac{(\omega_{(i,j)}.-\overline{\omega})^{2}}{N}}$
[ according to the rules of variation]
Therefore, $SSDLM_{(i,j))}$ is inversely proportional to homogeneity.
In order to determine the threshold of homogeneity (lower bound), background sampling was performed from a large population of images (10000 images). Random background pixels were then selected, and using a fixed kernel size $(5)$, the SSDLM was calculated. The mean and standard deviation (SD) of SSDLM were computed, and the threshold was set to $3\times$ SD in the positive scale, which was used as the SSDLM (homogeneity) threshold (lower bound) for a given modality. For bright-field microscopy, the homogeneity threshold was set to $4.23$. This value may vary depending on the modality or surface; for example, for phase-contrast microscopy, it is $3.5$, and for fiber sheets, it is approximately in the same range. However, this value can be adjusted by the user through tuning the corresponding hyperparameter.
2.3 Other Spatial Statistical Metric Used
2.3.1 Moran’s I
Moran’s I is a measure of spatial autocorrelation, which is defined as:
$I=\frac{n\sum_{i}\sum_{j}w_{ij}(x_{i}-\bar{x})(x_{j}-\bar{x})}{\sum_{i}\sum_{j}w_{ij}\cdot\sum_{i}(x_{i}-\bar{x})^{2}}$
$n:$ Number of spatial units, $x_{i}$, $x_{j}$: Values at locations i and j, where, $i$ and $j$ are the labels of two locations, $\bar{x}$: Mean of all values, $w_{ij}$: Spatial weight between locations i and j (e.g., inverse distance). $I>0$: Positive autocorrelation (clustering), $I<0$: Negative autocorrelation (dispersion) and $I=0$: No autocorrelation (randomness)
2.3.2 Cumulative Squared Shift of Nodal Intensity (CSSNI)
CSSNI quantifies local intensity variations in greyscale images by computing the sum of squared intensity differences (SSID) between each pixel and its 8-connected neighbour. The mathematical formulation of this metric is given below:
Let $\omega_{(i,j)}$ denote the intensity value of a pixel at the location $(i,j)$ in a patch size of $(M\times N)$ in a greyscale image. The CSSNI measure is defined as:
$CSSNI=\frac{1}{2}\sum_{i=1}^{M}\sum_{j=1}^{N}\sum_{(m,n)\in\mathcal{N}_{(i,j)}}(\omega_{(i,j)}-\omega_{(m,n)})^{2}$ (4)
where $\mathcal{N}_{(i,j)}$ represents the set of 8-connected neighbouring pixels. The division by 2 prevents double-counting since each pixel pair contributes to the sum twice.
2.3.3 Adjusted Variogram
Adjusted variogram is a variogram-like measure for spatial intensity variation. The spatial variation of the intensities of the pixels within an image patch is commonly analyzed using the semivariogram [15], which quantifies the relationship between the differences in the intensity of the pixels and their spatial separation. A conventional empirical semivariogram $\gamma(h)$ is defined as:
$\gamma(h)=\frac{1}{2N(h)}\sum_{i=1}^{N(h)}(x_{i}-x_{j})^{2}$
Where $h$ is the spatial lag, $x_{i}$ and $x_{j}$ are there pixel intensities at locations $i$ and $j$ (which are two different labels of the location), and $N(h)$ is the number of pixel pairs separated by $h$. Standard variograms rely on binning pixel pairs based on discrete lags to estimate spatial dependence.
This study introduces adjusted variogram providing a global estimate of intensity variation without explicit lag binning. The function computes the average squared intensity differences between all pixel pairs, normalised by their Euclidean distance in a specified neighbourhood. It can be formulated as:
$\gamma(h)=\frac{1}{2}\cdot\mathbb{E}\left[\frac{(z(x)-z(x+h))^{2}}{d(x,x+h)}\right]$ (5)
Where:
  • 1.
    $z(x)$ is the pixel intensity at position $x$,
  • 2.
    $z(x+h)$ is the pixel intensity at position $x+h$,
  • 3.
    $d(x,x+h)$ is the Euclidean distance between positions $x$ and $x+h$,
  • 4.
    $h$ is the spatial distance between pixels.
This method differs from conventional variograms in three key aspects:
  1. 1.
    No explicit lag binning: Instead of computing semivariance for specific lag distances, this approach aggregates all pairwise intensity differences.
  2. 2.
    Global averaging: The function provides a single scalar estimate of spatial intensity variation rather than a curve over multiple lags.
  3. 3.
    1D transformation: The patch is flattened into a sequence, which simplifies computation but does not fully preserve 2D spatial relationships.
The adjusted measure does not replace a traditional variogram; it serves as a computationally efficient alternative for capturing global intensity variation, here used with a goal of texture analysis and spatial feature extraction. Its overall advantages are i) fast and lightweight, ii) single-scalar output and iii) good for local analysis.
2.4 Computational Complexity Analysis
In this section, we analyze the computational complexity of the proposed methods to determine their efficiency in processing microscopic images.
2.4.1 Time Complexity of Functions
We analyze the time complexity of each function used in the algorithm. Here we denote the total number of pixels in the image as $N=n\times m$ and the patch size as $N_{p}$ (typically 5x5, 7x7 or 9x9).
  • 1.
    Calculating Moran’s I for a neighbourhood): The dominant terms arise from the distance matrix computation, weight matrix construction and dilation. Hence, the overall time complexity of the function is: $\mathcal{O}(N^{2})=\mathcal{O}((n\times m)^{2})$
  • 2.
    Calculating Adjusted Variogram-like Measure: This function computes pairwise distances and squared differences in $O(N_{p}^{2})$, leading to an overall complexity of $O(N_{p}^{2})$, where $N_{p}$ is the number of pixels in the patch.
  • 3.
    Calculating Cumulative Squared Shift of Nodal Intensity (CSSNI): This iterates over all pixels and their 8-connected neighbour, resulting in a time complexity of $O(N)$.
  • 4.
    Fuzzy Module: This function performs basic arithmetic and comparisons, leading to a constant time complexity of $O(1)$.
2.4.2 Overall Time Complexity
The most computationally expensive functions are calculating adjusted variogram and calculating Moran’s I, which run in $O(N_{p}^{2})$ per pixel ($p\leq 11$). Since these functions are not applied to every pixel in the image (used only where fuzzy cannot identify the pixels for the foreground or background set), the overall complexity of the algorithm will be much less than: $O(N\cdot N_{p}^{2}).$ Given that, $N_{p}$ is relatively small, the algorithm remains efficient for large images.
This analysis confirms that the proposed method is computationally feasible for high-resolution microscopic images while maintaining accuracy in segmentation and classification. However, Moran’s I can be accelerated using quadtree-based spatial indexing, Variogram calculations can be optimized with multi-resolution approximations. Also, parallel processing (Graphical Processing Unit (GPU) acceleration) can significantly reduce runtime for large images.
3 Related Work
In this section, a comprehensive overview is captured based on the contemporary research that introduces unsupervised learning based on several algorithmic & computational themes. Some notable research is included to depict the general evolutions of unsupervised learning for image segmentation problems. Unlike the images derived from the macro imaging system (which indicates the standard imaging system) or the images derived from the remote sensing system, the microscopic images have very high spatial resolution. In some microscopy modalities, specific contrast and brightness (phase contrast, fluorescence) are also important. The imagery often adopts special types of noises and other challenges which are already discussed in the introduction, and these noises or challenges are not present in the images obtained from standard or remote sensing systems. However, this literature review is not only restricted within the scope of unsupervised learning in the field of microscopy images. It examines the overall recent advancements and accompanying gaps in unsupervised image segmentation techniques, as the methodologies developed for unsupervised learning in segmentation problems are limited.
3.1 Unsupervised Learning for Image Segmentation Based on CNN
Among the several studies on unsupervised learning for image segmentation, one notable method is the convolutional neural network (CNN)-based algorithm for unsupervised image segmentation [16]. This approach optimizes feature extraction and clustering functions jointly, predicting cluster labels through differentiable functions. A spatial continuity loss enhances segmentation quality, while batch normalization normalizes response maps. PASCAL VOC 2012 and BSD500 are the used datasets for the experiment. However, the paper exhibits some research gaps as it does not address real-time segmentation performance, especially in the case of low-contrast noisy dataset which can be comparable to the scenario of microscopy images. Additionally, the paper has not argued scientifically about the diversity of input data and about the scalability of the proposed method.
3.2 Bayesian Statistical Model
An old study proposes the Bayesian unsupervised satellite image segmentation method based on stochastic estimation maximization (SEM) algorithm over global methods like MAP or MPM [17]. This study assesses spectral and spatial context contributions to image parameters. However, a serious research gap is discovered as the dependence on initialization affects solution reliability. Also, in certain cases, the limited exploration of spatial context contributions is observed. In case of low signal-to-noise ratio scenarios, the model needs to be developed for a good result and it should be validated for various image parameters. Another study, Image Segmentation with Adaptive Spatial Regularisation (ASR), introduced a Bayesian computation methodology accompanied by Potts-Markov random fields (MRFs) [18]. The method marginalized regularization parameters and considered small-variance asymptotic analysis. Despite achieving comparable results to supervised approaches, it lacked exploration of alternative regularization techniques and scalability assessments. Also, Independent pixel consideration neglects neighbouring pixel influence.
3.3 Soft Computing Based Approaches
A research involving Evolutionary Algorithm-Based Fuzzy Clustering (EABFC) introduces an unsupervised fuzzy clustering approach for image segmentation, combining an evolutionary algorithm (EA) with fuzzy clustering to leverage both local and non-local spatial information [19]. The method employs a multi-objective evolutionary sampling strategy to optimize pixel selection while preserving image details, followed by label correction using entropy and spatial constraints. However, the study has two key limitations: (1) it lacks experimental validation on diverse datasets, raising concerns about generalizability, and (2) it does not support user-defined parameters, limiting customization for different segmentation tasks. A novel approach combining fuzzy logic with Markov random field (MRF) has been proposed for image segmentation [20]. This method develops an adaptive fuzzy inference system and utilises spatial constraints effectively. The approach is notable for implementing a new clique potential MRF function. Fuzzy logic has been widely used for unsupervised segmentation. Fuzzy Random Fields and Unsupervised Image Segmentation proposed a fuzzy statistical model incorporating Gibbs sampling and stochastic estimation maximization (SEM) methods [21]. The approach demonstrated robust segmentation by integrating fuzzy components into traditional statistical models. However, the study did not explore real-world applications extensively or compare with advanced segmentation techniques. Another notable study, Estimation of Fuzzy Gaussian Mixture and Unsupervised Statistical Image Segmentation, applied adaptive iterative conditional estimation (ICE) to improve segmentation efficiency [22]. The model generalized statistical fuzzy segmentation and adapted it to contextual settings using SEM, ICE, and Expectation-Maximization (EM) algorithms. Future work aims to integrate the approach with existing segmentation techniques.
3.4 Unsupervised Domain Adaptation for Microscopy Images
Panoptic Domain Adaptive Mask (PDAM), based on Domain Adaptive Mask R-CNN (DAM), offers a novel segmentation strategy but struggles with domain shift due to contextual discrepancies [23]. This approach works with R-CNN and uses cycleGAN with an auxiliary objects inpainting mechanism. The former is responsible for synthesising images alike to the target, while the later one is responsible for reinforcing the image construction. The method encounters a problem regarding domain shift due to contextual information discrepancies and exhibits inadequate adaptation in the feature level for large domain gaps. Another approach puts forward an encoder-decoder-based multi-task learning model to cluster pixels according to foreground, background and cell boundaries as unsupervised domain adaptation. This method requires further improvements in domain-regularising cost functions and performance metrics, as the Dice metric is insensitive to clustered cell isolation. Also, the article does not discuss fixing the issue of ad-hoc parameter estimation [24].
3.5 Hyperspectral Image Segmentation
In another approach, hyperspectral microscopy image segmentation combines both unsupervised deep learning (UHRED) for denoising and supervised deep learning (SHRED) for enhancement followed by K-means clustering and mean squared error for loss calculation [25]. The method uses the Adam optimizer for determining the parameters of the model. However, challenges pertaining to overlapping species classification and automation of saturated pixel identification remain unanswered.
3.6 Classical and Hybrid Approaches
The traditional technique along with machine learning has tried to accomplish the unsupervised learning, competitive learning, fuzzy c-means clustering, and Gibbs random fields to improve tissue component segmentation through an iterative conditional modes (ICM) algorithm adaptation. The prominent lacks of this approach lie in limited accuracy produced, subjectivity in setting thresholds and the need for ICM algorithm adaptation [26]. Another research proposed a classical approach involving edge detection and morphological processes, called MPS-Based Image Segmentation for Bright-field Microscopy [27]. The method has been implemented in open-source software Fiji. Histogram equalisation, edge detection by the Canny edge detector and filling holes by using a maximum filter are the key steps to segment the total cell area by creating a binary image. Though it works with various cell types, it has a serious lack of handling the scenarios where the cell border is obscured. Also, the report about huge testing and comparison outcomes across diverse scenarios is absent. It requires better parameter standardisation across microscopy setups. Another work, Self-Supervised Learning (SSL 222 abbreviation replicated from publisher, ”Nature Communication”. https://www.nature.com/articles/s42003-025-08190-w) Approaches [31], advances annotation-free segmentation via optical flow-based pseudo-labeling, achieving $F_{1}$ scores of 0.77–0.88 on fluorescence images (as per the author’s claim). However, for bright-field myoblasts, it exhibits major limitations: (1) 50–60 s per image due to iterative optical flow, and (2) failure on 60% of low-contrast samples where texture features are unreliable. This hampers longitudinal studies needing speed and consistency. The error ”Either no cells found or all cells are touching the border” arises when cells are undetected or too close to image edges.
4 Methodology
The schematic diagram for end-to-end workflow is attached (Fig. 1). All the steps of the workflow are described in subsequent sections.
Figure 1: schematic diagram of end-to-end workflow
4.1 Data Acquisition
The first experiments used 10 primary real-time $1920\times 1440$ images (Dataset-1) of low-contrast myoblast cultures from our lab, featuring irregular, unstained cells with often obscured boundaries due to motion. The methodology was also tested on the LIVECell dataset (Dataset-2) [29], comprising diverse cell types and $3180$ cells. Finally, we segment the images (Dataset-3) of controllable laser traces on surface patterns and evaluate the robustness of our model across domains.
4.2 Colour Space Conversion
Most libraries (e.g., OpenCV) convert RGB to grayscale using a perceptual luminance model based on human vision sensitivity: $\text{Gray}=0.299R+0.587G+0.114B$. In contrast, this study applies an average-based conversion: $\text{Gray}=\frac{R+G+B}{3}$, assigning equal weights to all channels to ensure unbiased sensitivity in the treatment of second-order exponents of intensity variations across RGB components.
4.3 Primary Masking
In this step, the $SSDLM_{\omega}$ of the patches of all the pixels are checked. For a pixel and its patch, if its $SSDLM_{\omega}$ is found to be lower than the lower bound, the pixel is turned into black by replacing the pixel value with zero.
4.4 Fuzzy Inference System
Next, intensity transformation using fuzzy logic is performed, with sub-steps detailed below.
4.4.1 Fuzzy Membership Function Definitions
Three membership functions categorize pixel intensity, with parameters $\alpha=\beta=b=110$, $c=140$, and $a=80$ (Fig. 2).
Half-Trapezoidal Decreasing
$u_{d}(x)$
$u_{d}(I_{\text{eff}})=\begin{cases}1,&I_{\text{eff}}\leq a\\ \max\left(\frac{\alpha-I_{\text{eff}}}{\alpha-a},0\right),&\text{otherwise}\end{cases}$ (6)
Triangular
$u_{g}(x)$
$u_{g}(I_{\text{eff}})=\max\left(\min\left(\frac{I_{\text{eff}}-a}{b-a},\frac{c-I_{\text{eff}}}{c-b}\right),0\right)$ (7)
Figure 2: Membership functions $\mu_{dark}$, $\mu_{gray}$, and $\mu_{bright}$ for black, gray, and white regions.
Half-Trapezoidal Increasing
$u_{b}(x)$
$u_{b}(I_{\text{eff}})=\begin{cases}1,&I_{\text{eff}}\geq c\\ \max\left(\frac{I_{\text{eff}}-\beta}{c-\beta},0\right),&\text{otherwise}\end{cases}$ (8)
4.4.2 Hyperparameter Tuning
Default values are $\alpha=\beta=110$, $b=110$, $c=140$, $a=80$. These cannot be directly modified via the calibration window (Fig. 4), but the “Shift Gray” slider adjusts $b$ (default 110), shifting the gray midpoint. The “Span Gray” slider changes the slopes of the intersecting lines at $b$ without moving $b$ itself, thereby modifying $a$ and $c$ automatically by altering the $x$-coordinates where the lines meet the $X$-axis. As intensity is the independent variable, these adjustments apply to $x$-coordinates.
4.4.3 Fuzzy System
For each pixel, the grayscale intensity is evaluated using the membership functions to obtain $\mu_{dark}$, $\mu_{gray}$, and $\mu_{bright}$. The aggregated output is:
$I=\frac{v_{d}\,u_{d}(\text{input\_px})+v_{g}\,u_{g}(\text{input\_px})+v_{b}\,u_{b}(\text{input\_px})}{u_{d}(\text{input\_px})+u_{g}(\text{input\_px})+u_{b}(\text{input\_px})}$ (9)
where $v_{d}=0$, $v_{g}=127$, $v_{b}=255$. Based on rules, pixels with intensity <80<80 are black, >140>140 are white, and others are ambiguous, sent to the “Spatial Data Analysis” module.
4.5 Spatial Data Analysis
The ’Adjusted Variogram’ and ’Cumulative Squared Shift of Nodal Intensity’ (CSSNI) are computed and normalized by $SSDLM_{\omega}$ for each pixel based on its $5\times 5$ neighborhood. Pixels are first classified by the fuzzy system (intensity <80<80 as black, >140>140 as white), with ambiguous pixels subjected to spatial data analysis. Moran’s I ensures pixels in disordered neighborhoods are not misclassified as white, with this rule adjustable via the calibration window. While probabilistic approaches address pixel-level uncertainty [28], our method provides a complementary deterministic solution combining fuzzy logic and local spatial descriptors.
4.6 Final Masking
The final masking module resolves ambiguous pixels by analyzing intensity and spatial contrast in RGB channels. For mid-range fuzzy intensities, the normalized adjusted variogram within a $5\times 5$ neighborhood is evaluated: $v_{p}^{\text{norm}}=\frac{\gamma(\mathcal{N}_{p})}{\sigma(\mathcal{N}_{p})}$. Pixels below the ”NAV Threshold” (0–10, Fig. 4) undergo this analysis. Moran’s $I_{p}$ is computed over the same patch (”Randomness Threshold,” 1-1 to +1+1) to detect significant spatial structure, and a heuristic rule based on channel intensity and contrast is applied.
The rule considers green intensity $G_{p}$ relative to red and blue ($G_{p}<R_{p}$ and $G_{p}<B_{p}$), and the contrast ordering $d_{R}>d_{G}>d_{B}$, where $d_{*}$ is the normalized CSSNI per channel. Pixels meeting both conditions are classified as background; otherwise, they remain foreground.
Figure 3: Image after masking based on lower bound and upper bound where pink color denotes the uncertainty regions (left), Segmented image with Noise (Right)
4.7 Hyperparameter & Post-Segmentation Denoising:
Once the output is generated in the step called ”Final Masking”, Hyperparameter may be tuned for refining the output or denoising the output (if required) and replay the segmentation loop the desired result. Fig. 5 demonstrates two different example outputs generated by two types of tuning of hyperparameter for visual impact.
Figure 4: Left: Hyperparameter-Tuning Window, Right: Post-Segmentation Denoising Window
Post-Segmentation Denoising for (Dataset-1) begins with contour filling and morphological erosion of segmented objects (applicable for all datasets), followed by removing protein blobs based on circularity (isoperimetric quotient) and specified area thresholds, discarding blobs outside these bounds. In this regards, areas below 100 pixels are filled. Two circularity filters then remove objects: the first targets small objects (area 5–293, any circularity), and the second targets larger, irregular objects (area 253–1800, circularity << 0.31). A median blur (kernel size 5) is applied last (for all three datasets) to suppress salt-and-pepper noise. Fig. 4 illustrates the denoising GUI. The denoising profile varies with imaging modality and cell type; in this study, two profiles were used for 3,178 images:
Figure 5: Two different outputs generated by two hyperparameter tunings
Profile-1: Lower bound (LB) = 4.23 (default); pipeline: fill-below-threshold (area 100), erosion (kernel 3), circularity filter (area 0–71, circularity 0–1, mode=remove), median blur (kernel 5).
Profile-2: LB = 2.71; others are identical to Profile-1.
All images in the datasets 2 and 3 are processed with Profile-1, except for four categories of images in Dataset-2, where Profile-2 is used (Table 6).
4.8 Algorithm
This section includes two algorithms based on which the unsupervised masking works:
Algorithm 1: Unsupervised Masking
Input: RGB image $I$, patch size $n\times n$
Output: Binary mask $S$
01: Convert $I$ to grayscale: $I_{\text{gray}}=\frac{R+G+B}{3}$
02: Sample background pixels and compute SSDLM (Eq. 4)
03: Set threshold $L_{B}=\mu_{\text{SSDLM}}+3\sigma_{\text{SSDLM}}$
04: for each pixel $p\in I_{\text{gray}}$ do
05:  Extract patch $\mathcal{N}_{p}\in\mathbb{R}^{n\times n}$ around $p$
06:  Compute fuzzy value $f_{p}$ via fuzzy rules (Eqs. 7–9)
07:  if $f_{p}<a$ then
08:   Set $S(p)\leftarrow 0$
09:  else if $f_{p}>140$ then
10:   Set $S(p)\leftarrow 255$
11:  else /* ambiguous mid-range */
12:   Compute $v_{p}^{\text{norm}}=\gamma(\mathcal{N}_{p})/\sigma(\mathcal{N}_{p})$ (Eq. 6)
13:   if $v_{p}^{\text{norm}}<\texttt{NAV Threshold}$ then
14:    Compute Moran’s $I_{p}$ on $\mathcal{N}_{p}$
15:    if $I_{p}<\texttt{Randomness Threshold}$ then Set $S(p)\leftarrow 0$
16:    else
17:     Compute $d_{R},d_{G},d_{B}=\texttt{CSSNI}(\mathcal{N}_{R,G,B})$
18:     if $G_{p}<100\land G_{p}<R_{p}\land G_{p}<B_{p}$
19:      and $d_{R}>d_{G}>d_{B}$ then Set $S(p)\leftarrow 0$
20:     else Set $S(p)\leftarrow 255$
21:    end if
22:    Compute $d_{R},d_{G},d_{B}=\texttt{CSSNI}(\mathcal{N}_{R,G,B})$
23:    if $G_{p}<100\land G_{p}<R_{p}\land G_{p}<B_{p}$
24:     and $d_{R}>d_{G}>d_{B}$ then Set $S(p)\leftarrow 0$
25:    else Set $S(p)\leftarrow 255$
26:   end if
27:  end if
28: end for
29: if output $S$ requires refinement then
30:  Apply postprocessing (e.g., denoising)
31: else calibrate thresholds and repeat steps 03–29
4.9 Model Evaluation
To evaluate segmentation performance, we used a comprehensive set of standard metrics: Intersection over Union (IoU), pixel-wise accuracy, precision, recall and F1 score. Definitions and detailed formulations are standard in medical image analysis literature [30], and thus omitted here for brevity.
5 Results & Discussion
Result on Dataset-1:, the segmentation outcomes of 10 bright-field microscopy images of unstained myoblast (C2C12) cultures are presented.
Figure 6: Top-left: original image. Top-right: our model. Bottom-left: Cellpose. Bottom-right: StarDist. The green pixels are the true positives which are not detected by a particular model. The bright white pixels are true positives which are detected by a model too. The off-white pixels are false positives detected by a model
Table 1: Model Performance Comparison (Averages)
Model IoU Accuracy Precision Recall F1 Score
Our Model 0.431 0.871 0.531 0.726 0.601
StarDist 0.087 0.672 0.130 0.267 0.172
Cellpose 0.130 0.865 0.358 0.164 0.205
Table 2: Cohen’s Kappa Agreement Scores
Comparison Kappa ($\kappa$)
Expert 1 vs. Expert 2 0.81
Expert 1 vs. Our Model 0.78
Expert 2 vs. Our Model 0.75
Expert vs. Cellpose 0.42
Expert vs. StarDist 0.35
Each image was segmented using the proposed method, Cellpose, and StarDist for comparison. Ground truth masks were generated via expert annotations. Table 1 summarizes the average scores across 10 images. Table 4 represents the performance scores of the model with each of the 10 primary real-time images of live cells.
The proposed method outperforms both baselines in all metrics except for slightly lower accuracy compared to Cellpose, which is due to conservative masking (low false positive rate).
To validate qualitative segmentation consistency, two domain experts (cell biologists) independently assessed the outputs from all models on the original 10 images. Segmentations were rated as acceptable or not acceptable for morphological analysis. Cohen’s Kappa ($\kappa$) was computed to assess inter-observer reliability (Table 2).
The results indicate substantial agreement between the experts and the proposed method, supporting its reliability in low-contrast, label-free microscopy conditions.
Table 3: Wilcoxon Signed-Rank Test Results
textbfComparison Metric p-value
Our Model vs. Cellpose IoU 0.004
Our Model vs. StarDist IoU 0.002
Our Model vs. Cellpose F1-score 0.002
Our Model vs. StarDist F1-score 0.005
Figure 7: 3 more images + segmentations (ours vs. SOTA), ordered as Fig. 5. Full results for Dataset-1 are given in the location specified in abstract.
To statistically validate the superiority of the proposed method, the Wilcoxon signed-rank test (Table 3) was performed on per-image IoU and F1 scores. The tests compared our model’s results with Cellpose and StarDist across the same 10 images.
Table 4: Model Performance Comparison for 10 images
Img Model IoU Accuracy Precision Recall F1 Score
01 Our Model 0.39 0.83 0.54 0.58 0.56
01 Stardist 0.16 0.65 0.23 0.36 0.28
01 Cellpose 0.09 0.81 0.40 0.11 0.17
02 Our Model 0.43 0.95 0.50 0.75 0.60
02 Stardist 0.14 0.87 0.18 0.41 0.25
02 Cellpose <<.01 0.94 0.04 <<.01 <<.01
03 Our Model 0.47 0.88 0.57 0.74 0.64
03 Stardist 0.09 0.82 0.13 0.25 0.17
03 Cellpose 0.17 0.91 0.37 0.24 0.29
04 Our Model 0.36 0.96 0.50 0.56 0.52
04 Stardist 0.03 0.86 0.12 0.39 0.19
04 Cellpose 0.03 0.95 0.18 0.03 0.05
05 Our Model 0.43 0.89 0.53 0.70 0.61
05 Stardist 0.05 0.56 0.06 0.20 0.09
05 Cellpose 0.09 0.88 0.42 0.10 0.16
06 Our Model 0.38 0.82 0.40 0.87 0.55
06 Stardist 0.06 0.64 0.08 0.19 0.12
06 Cellpose 0.06 0.85 0.24 0.08 0.11
07 Our Model 0.39 0.82 0.40 0.91 0.56
07 Stardist 0.05 0.65 0.07 0.15 0.09
07 Cellpose 0.33 0.88 0.55 0.46 0.50
08 Our Model 0.43 0.87 0.50 0.77 0.61
08 Stardist 0.09 0.68 0.13 0.24 0.17
08 Cellpose 0.01 0.86 0.07 0.01 0.01
09 Our Model 0.55 0.91 0.64 0.79 0.71
09 Stardist 0.06 0.57 0.08 0.22 0.12
09 Cellpose 0.43 0.91 0.73 0.51 0.60
10 Our Model 0.48 0.78 0.73 0.59 0.65
10 Stardist 0.14 0.42 0.22 0.26 0.24
10 Cellpose 0.09 0.66 0.58 0.10 0.16
All p-values are below the 0.05 threshold, indicating that the improvements are statistically significant and not due to random chance.
For Dataset-1, visual analysis (Figure 6, 7) reveals that the proposed method handles low-contrast boundaries and uneven illumination more robustly than the baselines. StarDist and Cellpose fail in detecting shapes and show boundary leakage. Our method maintains structural fidelity in elongated or clustered cell regions by leveraging spatial statistics and fuzzy inference.
Table 5: Dataset-2: Segmentation performance metrics
Metric Mean Std
Dice 0.814431 0.078670
IoU 0.694019 0.106993
Accuracy 0.872015 0.096104
Precision 0.803413 0.112562
Recall 0.851245 0.124639
F1 Score 0.814431 0.078670
SSIM 0.502930 0.245851
Hausdorff 57.287395 34.920233
Figure 8: Two instances with segmented results taken from the Dataset-2 (LIVECell Public Dataset). (Left) Original image, (Right) Output. Full results for Dataset-2 are given in the location specified in abstract
Table 6: Image group statistics with mean F1 score and standard deviation for Profile-1
Image Group Count Mean F1 Std Profile
MCF7_Phase_F4 152 0.892839 0.046123 Profile-1
SkBr3_Phase_E3 151 0.883631 0.019668 Profile-1
SkBr3_Phase_H3 151 0.883416 0.023487 Profile-1
SkBr3_Phase_F3 146 0.879517 0.028173 Profile-1
MCF7_Phase_E4 157 0.874425 0.052069 Profile-1
MCF7_Phase_G4 160 0.866882 0.067715 Profile-1
A172_Phase_B7 128 0.845009 0.043283 Profile-1
A172_Phase_A7 129 0.842233 0.046283 Profile-1
A172_Phase_D7 128 0.826452 0.045234 Profile-1
BV2_Phase_D4 123 0.809159 0.066581 Profile-1
BT474_Phase_B3 147 0.808844 0.056174 Profile-1
BT474_Phase_C3 140 0.808419 0.057977 Profile-1
BV2_Phase_C4 128 0.805938 0.065642 Profile-1
BT474_Phase_A3 141 0.797301 0.063672 Profile-1
BV2_Phase_B4 133 0.795488 0.093175 Profile-1
SHSY5Y_Phase_D10 146 0.791682 0.040150 Profile-1
SHSY5Y_Phase_B10 156 0.788556 0.045465 Profile-1
SHSY5Y_Phase_C10 146 0.784836 0.044050 Profile-1
Huh7_Phase_A10 174 0.765936 0.069446 Profile-2
Huh7_Phase_A11 176 0.733752 0.096981 Profile-2
SKOV3_Phase_H4 139 0.720993 0.060473 Profile-2
SKOV3_Phase_G4 127 0.708126 0.077193 Profile-2
For Dataset-2 and Dataset-3 Fig 8 and Fig 9 are respectively attached as the representative images. Our method preserves structural fidelity in both elongated and clustered regions by combining spatial statistics with fuzzy logic-based inference. Table 6 shows group-wise average scores and segmentation profiles. While most categories used Profile-1, four of them used Profile-2. The analysis excluded eight corrupted mask instances. The average f1 scores for 3178 instances is 0.81 $\pm$ 0.07 (Tab  5).
The application of Profile-1 to the cell lines Huh7_Phase_A10, Huh7_Phase_A11, SKOV3_Phase_H4, and SKOV3_Phase_G4 results in the following mean F1-scores per category:
Huh7_Phase_A10: 0.6689(±0.0659)0.6689\ ($\pm$ 0.0659)
Huh7_Phase_A11: 0.6415(±0.0956)0.6415\ ($\pm$ 0.0956)
SKOV3_Phase_G4: 0.4742(±0.0817)0.4742\ ($\pm$ 0.0817)
SKOV3_Phase_H4: 0.5082(±0.0729)0.5082\ ($\pm$ 0.0729)
During the analysis, the following 8 images were excluded as their masks were found to be corrupted:
A172_Phase_A7_1_01d04h00m_3.png
A172_Phase_D7_1_01d20h00m_1.png
BT474_Phase_B3_1_03d00h00m_3.png
BV2_Phase_C4_1_01d16h00m_3.png
BV2_Phase_D4_1_00d12h00m_2.png
Huh7_Phase_A11_1_00d04h00m_3.png
Huh7_Phase_A11_1_00d04h00m_4.png
SHSY5Y_Phase_D10_1_01d16h00m_4.png
Figure 9: Segementation Result from Dataset-3 (Trace of laser); (left) Original Image, (Right) Output.
The combination of classical spatial statistics, fuzzy logic, and adaptive masking achieves strong unsupervised performance on a challenging modality. Both expert assessment and statistical testing confirm the model’s practical value. Hyperparameter tuning via GUI sliders enhances usability, and no pre-training is required, making it accessible for low-data environments.
6 Conclusion
This method enables unsupervised, training-free segmentation of bright-field microscopy images and is validated through standard performance metrics (Accuracy, Precision, Recall, IoU, F1-score), statistical significance testing, and expert visual assessment. While it allows calibration through an interactive interface, the segmentation process is largely automated and does not require annotated training data. This makes it well-suited for regenerative medicine applications, such as stem cell tracking, wound healing assays, and time-lapse live-cell imaging. The proposed approach supports fast, reproducible analysis in experimental and clinical microscopy workflows. Future directions encompass better precision and instance segmentation for overlapping cellular communities. Also, despite the promising results, our study has limitations pertaining to the size of the dataset and relatively low precision which would be addressed in future work.
Acknowledgments
The authors would like to thank Svetlana Ulasevich and Andrei Buzykin for providing microscopy images and for fruitful discussions of the domain area.
Funding
The research was supported by ITMO University Research Projects in AI Initiative (RPAII) (project #640103, Development of methods for automated processing and analysis of optical and atomic force microscopy images using machine learning techniques).
Declaration of Competing Interest
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper [^3] This publication is a preprint and has not been peer-reviewed. It is made available under the selected license. Any unauthorized use or violation of the license terms is strictly prohibited and will be the responsibility of the individual or entity engaging in such use..
Ethics Statement
This study did not involve human participants or animals.
Author Contributions
S.D. designed the study, analyzed the data, implemented the model, drafted the manuscript, and supervised the work. P.Z. provided supervision and critical review of the manuscript. All authors read and approved the final version.

### v2

@@
-GUI Based Fuzzy Logic and Spatial Statistics for Unsupervised Microscopy Segmentation
+Semi-Unsupervised Microscopy Segmentation with Fuzzy Logic and Spatial Statistics for Cross-Domain Analysis Using a GUI
 Abstract
-Brightfield microscopy imaging of unstained live cells remains a persistent challenge due to low contrast, temporal changes in specimen phenotypes, irregular illumination, and the absence of training labels.
-While deep learning (DL) methods (e.g., Cellpose 3.0) achieve state-of-the-art (SOTA) performance, they require extensive labeled data and heavy computational resources, and they often fail under uneven illumination.
-We present the first unsupervised segmentation framework combining spatial standard deviation from local mean (SSDLM), fuzzy logic, adjusted variograms, Moran’s I, and cumulative squared shift of nodal intensity (CSSNI) to address these limitations.
-Unlike deep learning models, our approach requires no annotations or retraining and operates through a user-friendly GUI tailored for non-programming users.
-The robustness and generality were validated on three datasets, including cross-domain data.
-We benchmark our method against 2023–2024 SOTA models, including Cellpose 3.0 and StarDist, using a dataset of unstained myoblast images.
-Our method achieves a significant improvement in segmentation performance, with an IoU increase of up to 48% and statistically validated superiority ($p<0.01$, Wilcoxon signed-rank test).
-Expert evaluation from two biologists further supports the segmentation quality (Cohen’s $\kappa>0.75$).
-The proposed algorithm is lightweight, interpretable, and computationally efficient, offering a practical and effective alternative for cell segmentation in label-free microscopy.
-The code, the dataset and the results are available for reproducibility [^1] The article is under review process.
-Code or Dataset may be provided upon request..
+Brightfield microscopy of unstained live cells is challenging due to low contrast, dynamic morphology, uneven illumination, and lack of labels.
+Deep learning achieved SOTA performance on stained, high-contrast images but needs large labeled datasets, expensive hardware, and fails under uneven illumination.
+This study presents a low-cost, lightweight, annotation-free segmentation method by introducing one-time calibration-assisted unsupervised framework adaptable across imaging modalities and image type.
+The framework determines background via spatial standard deviation from the local mean.
+Uncertain pixels are resolved using fuzzy logic, cumulative squared shift of nodal intensity, statistical features, followed by post-segmentation denoising calibration which is saved as a profile for reuse until noise pattern or object type substantially change.
+The program runs as a script or graphical interface for non-programmers.
+The method was rigorously evaluated using IoU, F1-score, and other metrics, with statistical significance confirmed via Wilcoxon signed-rank tests.
+On unstained brightfield myoblast (C2C12) images, it outperformed Cellpose 3.0 and StarDist, improving IoU by up to 48% (average IoU = 0.43, F1 = 0.60).
+In phase-contrast microscopy, it achieved a mean IoU of 0.69 and an F1-score of 0.81 on the LIVECell dataset ($n=3178$), with substantial expert agreement ($\kappa>0.75$) confirming cross-modality robustness.
+Successful segmentation of laser-affected polymer surfaces further confirmed cross-domain robustness.
+By introducing the Homogeneous Image Plane concept, this work provides a new theoretical foundation for training-free, annotation-free segmentation.
+The framework operates efficiently on CPU, avoids cell staining, and is practical for live-cell imaging and biomedical applications.
+Code and data are available for reproducibility.*
 keywords:
 Computer vision , image analysis , live cell imaging , quantitative microscopy , unsupervised segmentation
-journal: CMIG
+codecodefootnotetext: https://github.com/anonymusUID/01cvLiveCell01p
 1 Introduction
@@
 The study of biological samples, accompanied by machine learning and data science [2], often suffers from significant challenges for studying unstained (naturally pigmented) specimen which appear as low-contrast & noisy objects under microscope.
-Deep learning (DL) tools (e.g.Cellpose 3.0) [3] achieve impressive results for stained (high contrast or pigmented) labeled data, but they consistently fail under the low-contrast conditions characteristic of unstained samples due to their dependency on training data and sensitivity to noise.In addition, scarcity of sufficient annotated data hinders the DL training process.
+Deep learning (DL) tools (e.g.Cellpose 3.0) [3] achieve impressive results for stained (high contrast or pigmented) labeled data, but they consistently fail under the low-contrast conditions characteristic of unstained samples due to their dependency on training data and sensitivity to noise.
+In addition, scarcity of sufficient annotated data hinders the DL training process.
 Therefore, segmenting unstained live cells remains a significant challenge involving faster, annotation-free and training-free analysis of live cells without killing, staining, or fixing them.
@@
 (III) Optical aberrations – Dust, air bubbles, glare, and slide imperfections introduce artifacts.
-(IV) Overlapping structures – Cells and tissues often overlap which complicates segmentation.
+(IV) Overlapping dynamic structures – Cells and tissues often overlap with temporal changes in shape which complicates segmentation.
 (V) Annotation dependency – Microscopy data demands accurate annotation, requires expert knowledge, and is both error-prone and time-consuming.
@@
 Our Approach
-We present a fundamentally new framework that is found to be robust under the following assumptions: I) the environment involving optical conditions and the imaging system of bright-field microscopy is set up to achieve a nearly homogeneous image plane (defined in section §II.B); II) one of two image sets (either background or specimens) must have optical heterogeneity caused by its inherent texture.
+We present a fundamentally new framework that is found to be robust under the following assumptions: I) the environment involving optical conditions and the imaging system of bright-field microscopy is set up to achieve a nearly homogeneous image plane (defined in section 2); II) one of two image sets (either background or specimens) must have optical heterogeneity caused by its inherent texture.
 The propsed model combines:
@@
   • 1.
-    Unsupervised segmentation with ethical advantage: Our method eliminates the need for staining or labeling, preserving cell viability and avoiding cytotoxic effects—critical for live-cell and longitudinal studies.
+    Unsupervised segmentation for live cells: Our method eliminates the need for staining or labeling, preserving cell viability and avoiding cytotoxic effects—critical for live-cell and longitudinal studies.
   • 2.
@@
     Real-world validation: The framework is validated on challenging datasets, including primary cell cultures and clinical microscopy samples, demonstrating robust performance in diverse imaging conditions.
-2 Academic Discussion & Background
+2 Related Work
+In this section, a comprehensive overview is captured based on the contemporary research that introduces unsupervised learning based on several algorithmic & computational themes.
+Some notable research is included to depict the general evolutions of unsupervised learning for image segmentation problems.
+Unlike the images derived from the macro imaging system (which indicates the standard imaging system) or the images derived from the remote sensing system, the microscopic images have very high spatial resolution.
+In some microscopy modalities, specific contrast and brightness (phase contrast, fluorescence) are also important.
+The imagery often adopts special types of noises and other challenges which are already discussed in the introduction, and these noises or challenges are not present in the images obtained from standard or remote sensing systems.
+However, this literature review is not only restricted within the scope of unsupervised learning in the field of microscopy images.
+It examines the overall recent advancements and accompanying gaps in unsupervised image segmentation techniques, as the methodologies developed for unsupervised learning in segmentation problems are limited.
+2.1 Unsupervised Learning for Image Segmentation Based on CNN
+Among the several studies on unsupervised learning for image segmentation, one notable method is the convolutional neural network (CNN)-based algorithm for unsupervised image segmentation [16].
+This approach optimizes feature extraction and clustering functions jointly, predicting cluster labels through differentiable functions.
+A spatial continuity loss enhances segmentation quality, while batch normalization normalizes response maps.
+PASCAL VOC 2012 and BSD500 are the used datasets for the experiment.
+However, the paper exhibits some research gaps as it does not address real-time segmentation performance, especially in the case of low-contrast noisy dataset which can be comparable to the scenario of microscopy images.
+Additionally, the paper has not argued scientifically about the diversity of input data and about the scalability of the proposed method.
+2.2 Bayesian Statistical Model
+An old study proposes the Bayesian unsupervised satellite image segmentation method based on stochastic estimation maximization (SEM) algorithm over global methods like MAP or MPM [17].
+This study assesses spectral and spatial context contributions to image parameters.
+However, a serious research gap is discovered as the dependence on initialization affects solution reliability.
+Also, in certain cases, the limited exploration of spatial context contributions is observed.
+In case of low signal-to-noise ratio scenarios, the model needs to be developed for a good result and it should be validated for various image parameters.
+Another study, Image Segmentation with Adaptive Spatial Regularisation (ASR), introduced a Bayesian computation methodology accompanied by Potts-Markov random fields (MRFs) [18].
+The method marginalized regularization parameters and considered small-variance asymptotic analysis.
+Despite achieving comparable results to supervised approaches, it lacked exploration of alternative regularization techniques and scalability assessments.
+Also, Independent pixel consideration neglects neighbouring pixel influence.
+2.3 Soft Computing Based Approaches
+A research involving Evolutionary Algorithm-Based Fuzzy Clustering (EABFC) introduces an unsupervised fuzzy clustering approach for image segmentation, combining an evolutionary algorithm (EA) with fuzzy clustering to leverage both local and non-local spatial information [19].
+The method employs a multi-objective evolutionary sampling strategy to optimize pixel selection while preserving image details, followed by label correction using entropy and spatial constraints.
+However, the study has two key limitations: (1) it lacks experimental validation on diverse datasets, raising concerns about generalizability, and (2) it does not support user-defined parameters, limiting customization for different segmentation tasks.
+A novel approach combining fuzzy logic with Markov random field (MRF) has been proposed for image segmentation [20].
+This method develops an adaptive fuzzy inference system and utilises spatial constraints effectively.
+The approach is notable for implementing a new clique potential MRF function.
+Fuzzy logic has been widely used for unsupervised segmentation.
+Fuzzy Random Fields and Unsupervised Image Segmentation proposed a fuzzy statistical model incorporating Gibbs sampling and stochastic estimation maximization (SEM) methods [21].
+The approach demonstrated robust segmentation by integrating fuzzy components into traditional statistical models.
+However, the study did not explore real-world applications extensively or compare with advanced segmentation techniques.
+Another notable study, Estimation of Fuzzy Gaussian Mixture and Unsupervised Statistical Image Segmentation, applied adaptive iterative conditional estimation (ICE) to improve segmentation efficiency [22].
+The model generalized statistical fuzzy segmentation and adapted it to contextual settings using SEM, ICE, and Expectation-Maximization (EM) algorithms.
+Future work aims to integrate the approach with existing segmentation techniques.
+2.4 Unsupervised Domain Adaptation for Microscopy Images
+Panoptic Domain Adaptive Mask (PDAM), based on Domain Adaptive Mask R-CNN (DAM), offers a novel segmentation strategy but struggles with domain shift due to contextual discrepancies [23].
+This approach works with R-CNN and uses cycleGAN with an auxiliary objects inpainting mechanism.
+The former is responsible for synthesising images alike to the target, while the later one is responsible for reinforcing the image construction.
+The method encounters a problem regarding domain shift due to contextual information discrepancies and exhibits inadequate adaptation in the feature level for large domain gaps.
+Another approach puts forward an encoder-decoder-based multi-task learning model to cluster pixels according to foreground, background and cell boundaries as unsupervised domain adaptation.
+This method requires further improvements in domain-regularising cost functions and performance metrics, as the Dice metric is insensitive to clustered cell isolation.
+Also, the article does not discuss fixing the issue of ad-hoc parameter estimation [24].
+2.5 Hyperspectral Image Segmentation
+In another approach, hyperspectral microscopy image segmentation combines both unsupervised deep learning (UHRED) for denoising and supervised deep learning (SHRED) for enhancement followed by K-means clustering and mean squared error for loss calculation [25].
+The method uses the Adam optimizer for determining the parameters of the model.
+However, challenges pertaining to overlapping species classification and automation of saturated pixel identification remain unanswered.
+2.6 Classical and Hybrid Approaches
+The traditional technique along with machine learning has tried to accomplish the unsupervised learning, competitive learning, fuzzy c-means clustering, and Gibbs random fields to improve tissue component segmentation through an iterative conditional modes (ICM) algorithm adaptation.
+The prominent lacks of this approach lie in limited accuracy produced, subjectivity in setting thresholds and the need for ICM algorithm adaptation [26].
+Another research proposed a classical approach involving edge detection and morphological processes, called MPS-Based Image Segmentation for Bright-field Microscopy [27].
+The method has been implemented in open-source software Fiji.
+Histogram equalisation, edge detection by the Canny edge detector and filling holes by using a maximum filter are the key steps to segment the total cell area by creating a binary image.
+Though it works with various cell types, it has a serious lack of handling the scenarios where the cell border is obscured.
+Also, the report about huge testing and comparison outcomes across diverse scenarios is absent.
+It requires better parameter standardisation across microscopy setups.
+Another work, Self-Supervised Learning (SSL 111abbreviation replicated from publisher, ”Nature Communication”.
+https://www.nature.com/articles/s42003-025-08190-w) Approaches [31], advances annotation-free segmentation via optical flow-based pseudo-labeling, achieving $F_{1}$ scores of 0.77–0.88 on fluorescence images (as per the author’s claim).
+However, for bright-field myoblasts, it exhibits major limitations: (1) 50–60 s per image due to iterative optical flow, and (2) failure on 60% of low-contrast samples where texture features are unreliable.
+This hampers longitudinal studies needing speed and consistency.
+The error ”Either no cells found or all cells are touching the border” arises when cells are undetected or too close to image edges.
+3 Methodological Background & Technical Discussion
 3.1 Introducing Image Plane
@@
 Also, parallel processing (Graphical Processing Unit (GPU) acceleration) can significantly reduce runtime for large images.
-3 Related Work
-In this section, a comprehensive overview is captured based on the contemporary research that introduces unsupervised learning based on several algorithmic & computational themes.
-Some notable research is included to depict the general evolutions of unsupervised learning for image segmentation problems.
-Unlike the images derived from the macro imaging system (which indicates the standard imaging system) or the images derived from the remote sensing system, the microscopic images have very high spatial resolution.
-In some microscopy modalities, specific contrast and brightness (phase contrast, fluorescence) are also important.
-The imagery often adopts special types of noises and other challenges which are already discussed in the introduction, and these noises or challenges are not present in the images obtained from standard or remote sensing systems.
-However, this literature review is not only restricted within the scope of unsupervised learning in the field of microscopy images.
-It examines the overall recent advancements and accompanying gaps in unsupervised image segmentation techniques, as the methodologies developed for unsupervised learning in segmentation problems are limited.
-3.1 Unsupervised Learning for Image Segmentation Based on CNN
-Among the several studies on unsupervised learning for image segmentation, one notable method is the convolutional neural network (CNN)-based algorithm for unsupervised image segmentation [16].
-This approach optimizes feature extraction and clustering functions jointly, predicting cluster labels through differentiable functions.
-A spatial continuity loss enhances segmentation quality, while batch normalization normalizes response maps.
-PASCAL VOC 2012 and BSD500 are the used datasets for the experiment.
-However, the paper exhibits some research gaps as it does not address real-time segmentation performance, especially in the case of low-contrast noisy dataset which can be comparable to the scenario of microscopy images.
-Additionally, the paper has not argued scientifically about the diversity of input data and about the scalability of the proposed method.
-3.2 Bayesian Statistical Model
-An old study proposes the Bayesian unsupervised satellite image segmentation method based on stochastic estimation maximization (SEM) algorithm over global methods like MAP or MPM [17].
-This study assesses spectral and spatial context contributions to image parameters.
-However, a serious research gap is discovered as the dependence on initialization affects solution reliability.
-Also, in certain cases, the limited exploration of spatial context contributions is observed.
-In case of low signal-to-noise ratio scenarios, the model needs to be developed for a good result and it should be validated for various image parameters.
-Another study, Image Segmentation with Adaptive Spatial Regularisation (ASR), introduced a Bayesian computation methodology accompanied by Potts-Markov random fields (MRFs) [18].
-The method marginalized regularization parameters and considered small-variance asymptotic analysis.
-Despite achieving comparable results to supervised approaches, it lacked exploration of alternative regularization techniques and scalability assessments.
-Also, Independent pixel consideration neglects neighbouring pixel influence.
-3.3 Soft Computing Based Approaches
-A research involving Evolutionary Algorithm-Based Fuzzy Clustering (EABFC) introduces an unsupervised fuzzy clustering approach for image segmentation, combining an evolutionary algorithm (EA) with fuzzy clustering to leverage both local and non-local spatial information [19].
-The method employs a multi-objective evolutionary sampling strategy to optimize pixel selection while preserving image details, followed by label correction using entropy and spatial constraints.
-However, the study has two key limitations: (1) it lacks experimental validation on diverse datasets, raising concerns about generalizability, and (2) it does not support user-defined parameters, limiting customization for different segmentation tasks.
-A novel approach combining fuzzy logic with Markov random field (MRF) has been proposed for image segmentation [20].
-This method develops an adaptive fuzzy inference system and utilises spatial constraints effectively.
-The approach is notable for implementing a new clique potential MRF function.
-Fuzzy logic has been widely used for unsupervised segmentation.
-Fuzzy Random Fields and Unsupervised Image Segmentation proposed a fuzzy statistical model incorporating Gibbs sampling and stochastic estimation maximization (SEM) methods [21].
-The approach demonstrated robust segmentation by integrating fuzzy components into traditional statistical models.
-However, the study did not explore real-world applications extensively or compare with advanced segmentation techniques.
-Another notable study, Estimation of Fuzzy Gaussian Mixture and Unsupervised Statistical Image Segmentation, applied adaptive iterative conditional estimation (ICE) to improve segmentation efficiency [22].
-The model generalized statistical fuzzy segmentation and adapted it to contextual settings using SEM, ICE, and Expectation-Maximization (EM) algorithms.
-Future work aims to integrate the approach with existing segmentation techniques.
-3.4 Unsupervised Domain Adaptation for Microscopy Images
-Panoptic Domain Adaptive Mask (PDAM), based on Domain Adaptive Mask R-CNN (DAM), offers a novel segmentation strategy but struggles with domain shift due to contextual discrepancies [23].
-This approach works with R-CNN and uses cycleGAN with an auxiliary objects inpainting mechanism.
-The former is responsible for synthesising images alike to the target, while the later one is responsible for reinforcing the image construction.
-The method encounters a problem regarding domain shift due to contextual information discrepancies and exhibits inadequate adaptation in the feature level for large domain gaps.
-Another approach puts forward an encoder-decoder-based multi-task learning model to cluster pixels according to foreground, background and cell boundaries as unsupervised domain adaptation.
-This method requires further improvements in domain-regularising cost functions and performance metrics, as the Dice metric is insensitive to clustered cell isolation.
-Also, the article does not discuss fixing the issue of ad-hoc parameter estimation [24].
-3.5 Hyperspectral Image Segmentation
-In another approach, hyperspectral microscopy image segmentation combines both unsupervised deep learning (UHRED) for denoising and supervised deep learning (SHRED) for enhancement followed by K-means clustering and mean squared error for loss calculation [25].
-The method uses the Adam optimizer for determining the parameters of the model.
-However, challenges pertaining to overlapping species classification and automation of saturated pixel identification remain unanswered.
-3.6 Classical and Hybrid Approaches
-The traditional technique along with machine learning has tried to accomplish the unsupervised learning, competitive learning, fuzzy c-means clustering, and Gibbs random fields to improve tissue component segmentation through an iterative conditional modes (ICM) algorithm adaptation.
-The prominent lacks of this approach lie in limited accuracy produced, subjectivity in setting thresholds and the need for ICM algorithm adaptation [26].
-Another research proposed a classical approach involving edge detection and morphological processes, called MPS-Based Image Segmentation for Bright-field Microscopy [27].
-The method has been implemented in open-source software Fiji.
-Histogram equalisation, edge detection by the Canny edge detector and filling holes by using a maximum filter are the key steps to segment the total cell area by creating a binary image.
-Though it works with various cell types, it has a serious lack of handling the scenarios where the cell border is obscured.
-Also, the report about huge testing and comparison outcomes across diverse scenarios is absent.
-It requires better parameter standardisation across microscopy setups.
-Another work, Self-Supervised Learning (SSL 222 abbreviation replicated from publisher, ”Nature Communication”.
-https://www.nature.com/articles/s42003-025-08190-w) Approaches [31], advances annotation-free segmentation via optical flow-based pseudo-labeling, achieving $F_{1}$ scores of 0.77–0.88 on fluorescence images (as per the author’s claim).
-However, for bright-field myoblasts, it exhibits major limitations: (1) 50–60 s per image due to iterative optical flow, and (2) failure on 60% of low-contrast samples where texture features are unreliable.
-This hampers longitudinal studies needing speed and consistency.
-The error ”Either no cells found or all cells are touching the border” arises when cells are undetected or too close to image edges.
 4 Methodology
@@
 4.1 Data Acquisition
-The first experiments used 10 primary real-time $1920\times 1440$ images (Dataset-1) of low-contrast myoblast cultures from our lab, featuring irregular, unstained cells with often obscured boundaries due to motion.
-The methodology was also tested on the LIVECell dataset (Dataset-2) [29], comprising diverse cell types and $3180$ cells.
-Finally, we segment the images (Dataset-3) of controllable laser traces on surface patterns and evaluate the robustness of our model across domains.
+The first experiments used 10 primary real-time $1920\times 1440$ images (Dataset-1222Dataset avilable on github link.) of low-contrast myoblast (C2C12 cell) cultures from our lab, featuring irregular, unstained cells with often obscured boundaries due to motion.
+The cells were phographed using bright-field microscope.
+The methodology was also tested on the publicly avaiable LIVECell dataset (Dataset-2333https://sartorius-research.github.io/LIVECell/) [29], which comprises diverse cell types and 3,180 cells, with a modality of phase-contrast microscopy.
+Finally, we segment the images (Dataset-3444Dataset avilable on github link.) of controllable laser traces on surface patterns and evaluate the robustness of our model across domains.
 4.2 Colour Space Conversion
@@
 4.3 Primary Masking
-In this step, the $SSDLM_{\omega}$ of the patches of all the pixels are checked.
+In this step, the image background is considered a nearly homogeneous, monochromatic medium, and the $SSDLM_{\omega}$ of the patches of all the pixels are checked.
+This assumption is grounded in the lighting mechanism of Bright-Field microscopy; however, errors / uncertainties arising from several factors will be handled in subsequent steps.
 For a pixel and its patch, if its $SSDLM_{\omega}$ is found to be lower than the lower bound, the pixel is turned into black by replacing the pixel value with zero.
@@
 Definitions and detailed formulations are standard in medical image analysis literature [30], and thus omitted here for brevity.
-5 Results & Discussion
-Result on Dataset-1:, the segmentation outcomes of 10 bright-field microscopy images of unstained myoblast (C2C12) cultures are presented.
-Figure 6: Top-left: original image.
-Top-right: our model.
-Bottom-left: Cellpose.
-Bottom-right: StarDist.
-The green pixels are the true positives which are not detected by a particular model.
-The bright white pixels are true positives which are detected by a model too.
-The off-white pixels are false positives detected by a model
+4.10 Statistical Analysis
+Performance differences between the proposed framework and baseline models were evaluated using non-parametric statistical testing.
+The Wilcoxon signed-rank test was applied to per-image Intersection over Union (IoU) and F1-scores to determine whether observed improvements were statistically significant, with a significance threshold of $p<0.05$.
+To assess reliability of expert evaluation, inter-rater agreement between biologists and between each model and expert annotations was quantified using Cohen’s Kappa coefficient ($\kappa$), where values above 0.75 indicate substantial agreement.
+All statistical analyses were conducted at the per-image level to ensure fair and consistent comparisons across datasets.
+5 Results
+5.1 Dataset-1 (unstained myoblast C2C12 cell images, n = 10, with bright-field modality. The dimension of each image is $1920\times 1440$.)
 Table 1: Model Performance Comparison (Averages)
@@
 Cellpose 0.130 0.865 0.358 0.164 0.205
-Table 2: Cohen’s Kappa Agreement Scores
-Comparison Kappa ($\kappa$)
-Expert 1 vs. Expert 2 0.81
-Expert 1 vs. Our Model 0.78
-Expert 2 vs. Our Model 0.75
-Expert vs. Cellpose 0.42
-Expert vs. StarDist 0.35
-Each image was segmented using the proposed method, Cellpose, and StarDist for comparison.
-Ground truth masks were generated via expert annotations.
-Table 1 summarizes the average scores across 10 images.
-Table 4 represents the performance scores of the model with each of the 10 primary real-time images of live cells.
-The proposed method outperforms both baselines in all metrics except for slightly lower accuracy compared to Cellpose, which is due to conservative masking (low false positive rate).
-To validate qualitative segmentation consistency, two domain experts (cell biologists) independently assessed the outputs from all models on the original 10 images.
-Segmentations were rated as acceptable or not acceptable for morphological analysis.
-Cohen’s Kappa ($\kappa$) was computed to assess inter-observer reliability (Table 2).
-The results indicate substantial agreement between the experts and the proposed method, supporting its reliability in low-contrast, label-free microscopy conditions.
-Table 3: Wilcoxon Signed-Rank Test Results
-textbfComparison Metric p-value
-Our Model vs. Cellpose IoU 0.004
-Our Model vs. StarDist IoU 0.002
-Our Model vs. Cellpose F1-score 0.002
-Our Model vs. StarDist F1-score 0.005
-Figure 7: 3 more images + segmentations (ours vs. SOTA), ordered as Fig. 5. Full results for Dataset-1 are given in the location specified in abstract.
-To statistically validate the superiority of the proposed method, the Wilcoxon signed-rank test (Table 3) was performed on per-image IoU and F1 scores.
-The tests compared our model’s results with Cellpose and StarDist across the same 10 images.
 Table 2: Model Performance Comparison for 10 images
@@
 10 Cellpose 0.09 0.66 0.58 0.10 0.16
-All p-values are below the 0.05 threshold, indicating that the improvements are statistically significant and not due to random chance.
-For Dataset-1, visual analysis (Figure 6, 7) reveals that the proposed method handles low-contrast boundaries and uneven illumination more robustly than the baselines.
-StarDist and Cellpose fail in detecting shapes and show boundary leakage.
-Our method maintains structural fidelity in elongated or clustered cell regions by leveraging spatial statistics and fuzzy inference.
-Table 5: Dataset-2: Segmentation performance metrics
-Metric Mean Std
-Dice 0.814431 0.078670
-IoU 0.694019 0.106993
-Accuracy 0.872015 0.096104
-Precision 0.803413 0.112562
-Recall 0.851245 0.124639
-F1 Score 0.814431 0.078670
-SSIM 0.502930 0.245851
-Hausdorff 57.287395 34.920233
-Figure 8: Two instances with segmented results taken from the Dataset-2 (LIVECell Public Dataset).
-(Left) Original image, (Right) Output.
-Full results for Dataset-2 are given in the location specified in abstract
+The proposed method achieved an average Intersection over Union (IoU) of 0.431, accuracy of 0.871, precision of 0.531, recall of 0.726, and F1 score of 0.601 (Table 1).
+By comparison, Cellpose achieved IoU 0.130 and StarDist 0.087.
+Per-image performance is listed in Table 2.
+Statistical testing confirmed significance: the Wilcoxon signed-rank test (Table 3) showed all p-values << 0.01 when comparing our model against both Cellpose and StarDist.
+Expert evaluation indicated substantial agreement between biologists and the proposed model (Cohen’s $\kappa=0.781$ and $0.754$ vs experts, compared to $0.421$ for Cellpose and $0.351$ for StarDist; Table 4).
+Representative segmentations are shown in Figures 6 and 7.
+Figure 6: Top-left: original image.
+Top-right: our model.
+Bottom-left: Cellpose.
+Bottom-right: StarDist.
+The green pixels are the true positives which are not detected by a particular model.
+The bright white pixels are true positives which are detected by a model too.
+The off-white pixels are false positives detected by a model
+Figure 7: 3 more images + segmentations (ours vs. SOTA), ordered as Fig. 5. Full results for Dataset-1 are given in the location specified in abstract.
+Table 3: Wilcoxon Signed-Rank Test Results
+textbfComparison Metric p-value
+Our Model vs. Cellpose IoU 0.004
+Our Model vs. StarDist IoU 0.002
+Our Model vs. Cellpose F1-score 0.002
+Our Model vs. StarDist F1-score 0.005
+Table 4: Cohen’s Kappa Agreement Scores
+Comparison Kappa ($\kappa$)
+Expert 1 vs. Expert 2 0.81
+Expert 1 vs. Our Model 0.78
+Expert 2 vs. Our Model 0.75
+Expert vs. Cellpose 0.42
+Expert vs. StarDist 0.35
 Table 5: Image group statistics with mean F1 score and standard deviation for Profile-1
@@
 SKOV3_Phase_G4 127 0.708126 0.077193 Profile-2
-For Dataset-2 and Dataset-3 Fig 8 and Fig 9 are respectively attached as the representative images.
-Our method preserves structural fidelity in both elongated and clustered regions by combining spatial statistics with fuzzy logic-based inference.
-Table 6 shows group-wise average scores and segmentation profiles.
-While most categories used Profile-1, four of them used Profile-2.
-The analysis excluded eight corrupted mask instances.
-The average f1 scores for 3178 instances is 0.81 $\pm$ 0.07 (Tab  5).
-The application of Profile-1 to the cell lines Huh7_Phase_A10, Huh7_Phase_A11, SKOV3_Phase_H4, and SKOV3_Phase_G4 results in the following mean F1-scores per category:
-Huh7_Phase_A10: 0.6689(±0.0659)0.6689\ ($\pm$ 0.0659)
-Huh7_Phase_A11: 0.6415(±0.0956)0.6415\ ($\pm$ 0.0956)
-SKOV3_Phase_G4: 0.4742(±0.0817)0.4742\ ($\pm$ 0.0817)
-SKOV3_Phase_H4: 0.5082(±0.0729)0.5082\ ($\pm$ 0.0729)
-During the analysis, the following 8 images were excluded as their masks were found to be corrupted:
-A172_Phase_A7_1_01d04h00m_3.png
-A172_Phase_D7_1_01d20h00m_1.png
-BT474_Phase_B3_1_03d00h00m_3.png
-BV2_Phase_C4_1_01d16h00m_3.png
-BV2_Phase_D4_1_00d12h00m_2.png
-Huh7_Phase_A11_1_00d04h00m_3.png
-Huh7_Phase_A11_1_00d04h00m_4.png
-SHSY5Y_Phase_D10_1_01d16h00m_4.png
+5.2 Dataset-2 (LIVECell, n = 3178)
+The framework achieved mean Dice score of 0.8144 $\pm$ 0.0786, IoU of 0.6940 $\pm$ 0.1069, accuracy of 0.8720 $\pm$ 0.0961, precision of 0.8034 $\pm$ 0.1125, and recall of 0.8512 $\pm$ 0.1246 (Table 6).
+Structural similarity (SSIM) was 0.5029 $\pm$ 0.2458, and average Hausdorff distance was 57.2873 $\pm$ 34.9202.
+Group-wise statistics are reported in Table 5.
+Most cell categories performed with Profile-1, while four categories (Huh7_Phase_A10, Huh7_Phase_A11, SKOV3_Phase_G4, SKOV3_Phase_H4) required Profile-2, yielding mean F1 scores of 0.6689 $\pm$ 0.0659, 0.6415 $\pm$ 0.0956, 0.4742 $\pm$ 0.0817, and 0.5082 $\pm$ 0.0729 respectively.
+Representative results are shown in Figure 8.
+During the analysis, the following eight images were excluded as their masks were found to be corrupted: A172_Phase_A7_1_01d04h00m_3.png, A172_Phase_D7_1_01d20h00m_1.png, BT474_Phase_B3_1_03d00h00m_3.png, BV2_Phase_C4_1_01d16h00m_3.png, BV2_Phase_D4_1_00d12h00m_2.png, Huh7_Phase_A11_1_00d04h00m_3.png, Huh7_Phase_A11_1_00d04h00m_4.png, SHSY5Y_Phase_D10_1_01d16h00m_4.png
+Table 6: Dataset-2: Segmentation performance metrics
+Dice IoU Accuracy Precision Recall F1 Score SSIM Hausdorff
+Mean 0.814431 0.694019 0.872015 0.803413 0.851245 0.814431 0.502930 57.287395
+Std 0.078670 0.106993 0.096104 0.112562 0.124639 0.078670 0.245851 34.920233
+Figure 8: Two instances with segmented results taken from the Dataset-2 (LIVECell Public Dataset).
+(Left) Original image, (Right) Output.
+Full results for Dataset-2 are given in the location specified in abstract
+5.3 Dataset-3 (Controllable Laser Trace)
+The framework maintained consistent segmentation quality (with F1-score 0.830.900.83-0.90) on simulated laser-trace images, as shown in Figure 9.
 Figure 9: Segementation Result from Dataset-3 (Trace of laser); (left) Original Image, (Right) Output.
-The combination of classical spatial statistics, fuzzy logic, and adaptive masking achieves strong unsupervised performance on a challenging modality.
-Both expert assessment and statistical testing confirm the model’s practical value.
-Hyperparameter tuning via GUI sliders enhances usability, and no pre-training is required, making it accessible for low-data environments.
+6 Discussion
+The results demonstrate that the proposed framework consistently outperforms Cellpose and StarDist on unstained brightfield images, achieving up to 48% higher IoU on Dataset-1.
+All Wilcoxon signed-rank tests comparing the proposed method with Cellpose and StarDist yielded $p$-values below 0.05, confirming that the observed improvements in IoU and F1-scores were statistically significant rather than random variation.
+Furthermore, Cohen’s Kappa analysis indicated substantial agreement ($\kappa>0.75$) between expert annotations and the proposed framework, reinforcing the reliability of the segmentation outputs in practical biomedical applications related to hard challenge – low-contrast microscopy.
+The findings validate the central concept of the Homogeneous Image Plane.
+In microscopy, objects appear heterogeneous due to scattering and texture, while the background remains nearly homogeneous.
+By explicitly exploiting this property with spatial statistics and fuzzy inference, the framework separates object heterogeneity from background homogeneity.
+Compared with deep learning approaches, our method has three key advantages:
+  1. 1.
+    Training-free and annotation-free — no labeled data or retraining needed.
+  2. 2.
+    Robustness across modalities — three profiles sufficed for 3200+ images from diverse cell lines. Calibration is required only when imaging modality or specimen type changes substantially.
+  3. 3.
+    Accessibility — runs efficiently on CPU hardware, usable via both GUI / scripts, and suited to laboratories without GPUs or annotation resources.
+A limitation of the study is the relatively small size of Dataset-1.
+However, this was addressed by validation on the large LIVECell dataset ($n>3000$).
+Precision values were somewhat lower than Cellpose, reflecting a conservative design that prioritizes recall and structural fidelity — a trade-off beneficial in regenerative medicine, where false negatives are more critical than false positives.
+Future extensions include automating calibration profile selection, expanding validation to additional modalities, and integrating explainable AI for improved interpretability.
+Overall, the framework introduces a novel theoretical foundation through the Homogeneous Image Plane concept and demonstrates practical robustness across microscopy modalities, offering an alternative to data-hungry deep learning methods.
 7 Conclusion
-This method enables unsupervised, training-free segmentation of bright-field microscopy images and is validated through standard performance metrics (Accuracy, Precision, Recall, IoU, F1-score), statistical significance testing, and expert visual assessment.
+This method enables robust, unsupervised, training-free segmentation of bright-field microscopy images and is validated through standard performance metrics (Accuracy, Precision, Recall, IoU, F1-score), statistical significance testing, and expert visual assessment.
 While it allows calibration through an interactive interface, the segmentation process is largely automated and does not require annotated training data.